# TEncDM: Understanding the Properties of Diffusion Model in the Space of Language Model Encodings

### v2

@@ Abstract
This paper presents the Text Encoding Diffusion Model (TEncDM), a novel approach to diffusion modeling that operates in the space of pre-trained language model encodings. In contrast to traditionally used embeddings, encodings integrate contextual information. In our approach, we also employ a transformer-based decoder, specifically designed to incorporate context in the token prediction process. We conduct a comprehensive examination of the influence of the encoder, decoder, noise scheduler, and self-conditioning on zero-shot generation. Furthermore, we compare TEncDM with previous approaches on three conditional text generation tasks: QQP, XSum, and Wiki-Auto. The results show that TEncDM exhibits superior performance compared to existing non-autoregressive diffusion models.
@@ 2 Problem Statement and Background
The standard diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2021) learn to sample data from an unknown distribution by gradually denoising random Gaussian noise. The training procedure is defined through a forward diffusion process that satisfies $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\alpha_{t}}z_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}\in[0,1]$ is a predefined noise schedule, $t\in[0,1]$ and $\alpha_{t}>\alpha_{t+\Delta t}$. The denoising network (parameterized by $\theta$) is trained to reconstruct the original latent $z_{0}$ given the noisy latent $z_{t}$, as expressed in equation 1.
@@
Sampling procedure starts from a pure Gaussian noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ and utilizes the denoising network to iteratively generate latents $z_{t_{T-1}},...,z_{t_{1}}$, where $01=t_{T}>t_{T-1}>...>t_{1}=$ 0.
@@ 5.3 Diffusion model, $\hat{z}_{\theta}$
The diffusion model consists of 12 BERT layers and it is trained to reconstruct the original latent $z_{0}$ given its noisy version $z_{t}$ and a timestep $t$ by minimizing the objective (1). We provide the model with information about the timestep by adding its embedding to the hidden state vectors of each layer.
@@ 5.4 Generation process
For the conditional generation we keep the framework design similar to unconditional generation. The only difference is that we add condition encoder to process the input text and provide both diffusion model and decoder with its output via cross-attention. Implementation details can be found in Appendix E.
@@
Table 1: Comparison of diffusion encoders.
6 Datasets
7 Empirical Analysis
@@
In this section, we evaluate the components of our framework on the ROCStories and Wikipedia datasets. To simplify the setup, we only consider unconditional generation. In Section 8, we demonstrate that our findings can be successfully transferred to the conditional generation problems. In this section, we do not compare our method with others. The comparison with the GPT-2 (Radford et al. 2019) on unconditional generation is presented in Appendix I.
7.1 Evaluation Metrics
To calculate all the metrics, we generate 1000 texts. For MAUVE, we sample 1000 reference texts from the test set. We repeat this procedure 5 times and report the mean and standard deviation of the obtained results in $\text{mean}_{\text{std}}$ notation.
7.2 Model setup
The training of TEncDM is conducted within the latent space of BERT encodings, as it has shown the best performance among all encoders. We employ a 3-layer transformer for the decoder and train it to reconstruct $z_{0}$ from $z_{t}$, where $t\in U[0,0.15]$. A comprehensive analysis of various decoder modifications is presented in Section 7.2 and Appendix B. The diffusion model is the 12-layer transformer with dimensionality of 768. By default we train it with tan-9 noise scheduler.
Effect of Diffusion Encoder
We compare latent spaces of BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), BART (Lewis et al. 2020) and T5 (Raffel et al. 2020) encoders, as well as BERT embeddings, to ascertain the optimal choice for the diffusion model. All encoders have approximately the same size of 100M parameters. In this experiment, we train diffusion models with the same set of hyperparameters across all diffusion encoders. We train the decoders according to the scheme described in Section 7.2. The results of this comparison are presented in Table 1 and they show a clear advantage of the latent space derived from BERT encodings on ROCStories dataset. Furthermore, the quality of all encoders is superior to that of BERT embeddings. A better div and mem for embeddings can be explained by the presence of words in the corpus that do not align with the context. The text samples are presented Table 9 of Appendix J. This confirms our hypothesis that encodings are better suited for the training of a diffusion model. We discuss the drop in mauve for RoBERTa in Appendix G.
Effect of Decoder
Figure 2: Comparison of models with and without self-conditioning on ROCStories dataset.
Figure 3: Comparison of prediction magnitudes for generation processes with different amount of steps on ROCStories dataset.
@@ 7.2 Model setup
Figure 4: Comparison of noise schedulers on ROCStories dataset.
@@
Table 3: Comparison of noise schedulers.
@@
Table 4: Seq2Seq evaluation results of AR and Diffusion methods on QQP, XSum and Wiki-Auto datasets. We calculate ROUGE-1/2/L (R-1/2/L), BERTScore (BS) and BLEU-4 (B-4). Result of AR approaches, DiffuSeq, and SeqDiffuSeq were sources from their respective publications (marked as $\star$). Additionally, we trained AR-Diffusion and SeqDiffuSeq on previously unreported datasets, adhering to the methodologies outlined by the authors (marked as $\dagger$).
@@ 8 Seq2Seq Experiments
We include two groups of baselines in comparison. The first group comprises of classical AR baselines: Transformer (Vaswani et al. 2017a), FLAN-T5-base (Chung et al. 2024) and finetuned GPT-2-small (Radford et al. 2019). Besides, we compare TEncDM to popular embedding-based diffusion methods: DiffuSeq (Gong et al. 2023), SeqDiffuSeq (Yuan et al. 2024), GENIE (Lin et al. 2023), AR-diffusion (Wu et al. 2023). We focus only on non-autoregressive diffusion models trained in the latent space of embeddings or encodings.
@@
For evaluation of paraphrasing and simplification tasks, we adopt the setting of SeqDiffuSeq (Yuan et al. 2024) and calculate ROUGE-L (Lin 2004), BERTScore (Zhang et al. 2020) and BLEU-4. In addition, we follow the approach of Wu et al. (2023) and report ROUGE-1/2/L for summarization task.
@@ Results
Table 4 presents a comprehensive comparison of our approach against existing methods across all datasets. Results for AR approaches, DiffuSeq, and SeqDiffuSeq were sourced from their respective publications (Qi et al. 2021; Wu et al. 2023; Yuan et al. 2024). Furthermore, we trained both AR-Diffusion and SeqDiffuSeq on previously unreported datasets, adhering to the methodologies outlined by the authors.
We experiment with three encoders: BERT-base, T5-base, RoBERTa-base to investigate their efficacy on conditional tasks. We use the same encoder for $E_{diff}$ and $E_{cond}$. Our findings indicate that all three encoders demonstrate effectiveness across the tasks, achieving comparable performance levels. However, no single encoder consistently outperforms the others across all tasks. T5-base excels in question-question paraphrasing (QQP) across all metrics, while RoBERTa-base demonstrates superior performance on extreme summarization (XSum). BERT-base and T5-base exhibit comparable performance on text simplification.
A comparison against other methods clearly demonstrates that TEncDM, utilizing any of the tested encoder models, outperforms popular diffusion embedding-based approaches. Furthermore, TEncDM-T5 achieves either superior or equivalent performance to all classical AR baselines on QQP. Similarly, TEncDM-RoBERTa demonstrates comparable quality against FLAN-T5-base on extreme summarization (XSum). Furthermore, TEncDM surpasses diffusion and AR approaches by a large margin on simplification task.
@@ 9 Limitations
There are three limitations that warrant further investigation. First, while the quality of the model can be improved by training diffusion encoder, decoder and denoising model simultaneously, we avoid doing so in order to avoid overcomplicating the approach. Second, samples from the latent space have a high dimensionality that depends on the sequence length, making the training of our method significantly slower as the length increases. This problem can probably be eliminated by training the autoencoder, which is a great direction for the further research. Third, as different encoders works better for different tasks, it is necessary to find the best one for each task.
10 Conclusion
In this work, we explore key details of the diffusion pipeline for text generation. We propose TEncDM which trains the diffusion model in the latent space of language model encoding. In contrast to embeddings, they contain contextual information which helps diffusion model to recover latents. In order to improve text generation performance, we analyse the effect of self-conditioning and conclude that it increases the magnitudes of the model’s predictions and results in reducing of generation steps number. Additionally, we propose an efficient decoder and noise scheduler that boost the diffusion model performance. The extensive ablation on ROCStories and Wikipedia proves the impact of proposed design choices. Finally, TEncDM outperforms recent diffusion models and classical autoregressive methods in downstream task experiments.
@@ Appendix B Corruption for decoder training
    Replacing $z_{0}$ with $z_{t}$ by the diffusion forward process, $Cor(z_{0})=\sqrt{\alpha_{t}}z_{0}+\sqrt{(1-\alpha_{t})}\varepsilon=z_{t}$.
@@ C.2 Wikipedia self-conditioning analysis
Figure 9 depicts the trajectories of prediction magnitudes for varying amounts of generation steps in the Wikipedia dataset. It can be observed that the trajectories exhibit a similar trend to those observed in ROCStories, although the spread of final prediction magnitudes is larger.
@@ Appendix E Implementation details
Table 6: Training details for TEncDM across different datasets.
@@ Appendix G Low MAUVE for RoBERTa encoder
Appendix H Scalability
We investigate the scalability of TEncDM. Our model leverages both a text encoder and a diffusion model, offering two avenues for parameter expansion: scaling the diffusion model alone or scaling both the encoder and diffusion model concurrently. In this section, we focus on the first approach, where we exclusively increase the size of the diffusion model. Specifically, we augmented the number of transformer heads from 12 to 16 and doubled the number of layers from 12 to 24. We call the standard model “base” and the increased model “large”. In Table 8 we show comparison of base and large diffusion models on Wikipedia dataset. The results demonstrate the effectiveness of the scaling strategy. The second approach, involving simultaneous scaling of both the encoder and diffusion model, necessitates significantly greater computational resources and training time. Therefore, we defer the exploration of this approach to future research.
@@ Appendix H Scalability
Appendix I Comparison with GPT-2
@@
Table 10: Comparison on unconditional generation.
Appendix J Generation examples

### v3

# TEncDM: Understanding the Properties of the Diffusion Model in the Space of Language Model Encodings

@@ Abstract
This paper presents the Text Encoding Diffusion Model (TEncDM), a novel approach to diffusion modeling that operates in the space of pre-trained language model encodings. In contrast to traditionally used embeddings, encodings integrate contextual information. In our approach, we also employ a transformer-based decoder, specifically designed to incorporate context in the token prediction process. We conduct a comprehensive examination of the influence of the encoder, decoder, noise scheduler, and self-conditioning on zero-shot generation. Furthermore, we compare TEncDM with previous approaches on three conditional text generation tasks: QQP, XSum, and Wiki-Auto. The results show that TEncDM exhibits superior performance compared to existing non-autoregressive diffusion models. Our code is available at https://github.com/M0RJIQUE/tencdm.
@@ 2 Problem Statement and Background
Gaussian diffusion models (Song et al. 2021) learn to sample data from an unknown distribution by gradually denoising random Gaussian noise. The training procedure is defined through a forward diffusion process that satisfies $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\alpha_{t}}z_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}\in[0,1]$ is a predefined noise schedule, $t\in[0,1]$ and $\alpha_{t}>\alpha_{t+\Delta t}$. The denoising network (parameterized by $\theta$) is trained to reconstruct the original latent $z_{0}$ given the noisy latent $z_{t}$, as expressed in equation 1.
@@
Sampling procedure starts from a pure Gaussian noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ and utilizes the denoising network to iteratively generate latents $z_{t_{T-1}},...,z_{t_{1}}$, where $01=t_{T}>...>t_{1}=$ 0.
@@ 5.3 Diffusion model, $\hat{z}_{\theta}$
The diffusion model consists of 12 BERT layers and it is trained to predict the original latent $z_{0}$ given its noisy version $z_{t}$ and a timestep $t$ by minimizing the objective (1). We provide the model with timestep value by adding timestep embedding to the hidden state vectors on each layer.
@@ 5.4 Generation process
For the conditional generation we keep the framework design similar to unconditional generation. The only difference is that we add condition encoder to process the input text and provide both diffusion model and decoder with its output via cross-attention. We also add classifier-free guidance with coefficient 0.5, because it slightly improves quality. Implementation details can be found in Appendix E.
6 Datasets
7 Empirical Analysis
In this section, we evaluate the components of our framework on the ROCStories and Wikipedia datasets. To simplify the setup, we only consider unconditional generation. In Section 8, we demonstrate that our findings can be successfully transferred to the conditional generation problems. In this section, we do not compare our method with others. The comparison with the GPT-2 (Radford et al. 2019) on unconditional generation is presented in Appendix J.
7.1 Evaluation Metrics
To calculate all the metrics, we generate 1000 texts. For MAUVE, we sample 1000 reference texts from the test set. We repeat this procedure 5 times and report the mean and standard deviation of the results in the $\text{mean}_{\text{std}}$ notation.
7.2 Model setup
Unless otherwise stated, the training of TEncDM is performed within the latent space of BERT encodings. A three-layer transformer is employed for the decoder, which is trained to reconstruct $z_{0}$ from $z_{t}$, where $t\in U[0,0.15]$. A comprehensive analysis of various decoder modifications is presented in this section and Appendix B. The diffusion model is a 12-layer transformer with a dimensionality of 768. We train it with tan-9 noise scheduler.
Effect of Diffusion Encoder
We compare latent spaces of BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), BART (Lewis et al. 2020) and T5 (Raffel et al. 2020) encoders, as well as BERT embeddings, to ascertain the optimal choice for the diffusion model. All encoders have approximately the same size of 100M parameters. In this experiment, we train diffusion models with the same set of hyperparameters across all diffusion encoders. We train the decoders according to the scheme described in Section 7.2. The results of this comparison are presented in Table 1 and they show a clear advantage of the latent space derived from BERT encodings on ROCStories dataset. Furthermore, the quality of all encoders is superior to that of BERT embeddings. A better div and mem for embeddings can be explained by the presence of words in the corpus that do not align with the context. The text samples are presented Table 9 of Appendix K. This confirms our hypothesis that encodings are better suited for the training of a diffusion model. We discuss the drop in mauve for RoBERTa in Appendix G.
@@ 7.2 Model setup
Table 1: Quality of diffusion models trained with different diffusion encoders.
Effect of Decoder
@@
Figure 2: Generation quality of diffusion models with and without self-conditioning on ROCStories dataset.
Figure 3: Prediction magnitudes for generation processes with different amount of steps on ROCStories dataset.
@@
Figure 4: Reconstruction loss and reconstruction accuracy of diffusion models trained with different noise schedulers on ROCStories dataset.
@@
Table 3: Quality of diffusion models trained with different noise schedulers.
@@
Table 4: Seq2Seq evaluation results of Diffusion and AR methods on QQP, XSum and Wiki-Auto datasets. We calculate ROUGE-1/2/L (R-1/2/L), BERTScore (BS) and BLEU-4 (B-4). All results taken from other papers are marked with $\star$. DiffuSeq and SeqDiffuSeq results were taken from their respective publications. Results for AR models were taken from (Gong et al. 2023; Wu et al. 2023; Lovelace et al. 2023). Additionally, we trained AR-Diffusion and SeqDiffuSeq on previously unreported datasets, using the code from the corresponding papers (marked as $\dagger$).
@@ 8 Seq2Seq Experiments
We include two groups of baselines in comparison. The first group comprises of popular diffusion methods: DiffuSeq (Gong et al. 2023), SeqDiffuSeq (Yuan et al. 2024), GENIE (Lin et al. 2023), AR-diffusion (Wu et al. 2023). We focus only on non-autoregressive diffusion models trained in the latent space of embeddings or encodings. Besides, we compare TEncDM to classical AR baselines: Transformer (Vaswani et al. 2017a), FLAN-T5-base (Chung et al. 2024) and finetuned GPT-2-small (Radford et al. 2019).
@@
For evaluation of paraphrasing and simplification tasks, we adopt the setting of SeqDiffuSeq (Yuan et al. 2024) and calculate ROUGE-L (Lin 2004), BERTScore (Zhang et al. 2020) and BLEU-4. In addition, we follow the approach of (Wu et al. 2023) and report ROUGE-1/2/L for summarization task.
@@ Results
Table 4 presents a comprehensive comparison of our approach against existing methods across three datasets. Results for DiffuSeq, and SeqDiffuSeq were sourced from their respective papers (Wu et al. 2023; Yuan et al. 2024). Results for GENIE were taken from the (Wu et al. 2023), as the original paper used ground truth labels to select the best generated text, which introduces unfairness. Additionally, we trained both AR-Diffusion and SeqDiffuSeq on previously unreported datasets, using the code from the corresponding papers. Results for AR models were taken from (Gong et al. 2023; Wu et al. 2023; Lovelace et al. 2023).
We experiment with three encoders: BERT-base, T5-base, RoBERTa-base to investigate their efficacy on conditional tasks. We use the same encoder for $E_{diff}$ and $E_{cond}$. Our findings indicate that all three encoders demonstrate effectiveness across the tasks, achieving comparable performance levels. However, no single encoder outperforms the others across all tasks. T5-base excels in question paraphrasing (QQP), while RoBERTa-base demonstrates superior performance on summarization (XSum) and on text simplification (Wiki-Auto) all encoders exhibit similar performance.
A comparison with other methods clearly shows that TEncDM, using any of the tested encoder models, outperforms popular diffusion embedding-based approaches. Furthermore, TEncDM with an optimal encoder achieves comparable performance to three AR baselines on QQP and XSum and outperforms them on Wiki-Auto.
9 Conclusion
In this work, we explore key details of the diffusion pipeline for text generation. We propose TEncDM which trains the diffusion model in the latent space of the language model encoding. Unlike embeddings, they contain contextual information which helps diffusion model to recover latents. To improve text generation performance, we analyse the effect of self-conditioning and conclude that it increases the magnitude of the model’s predictions, which allows to reduce the number of generation steps. In addition, we propose an efficient decoder and noise scheduler that boost the generation quality. Extensive ablation on ROCStories and Wikipedia datasets demonstrates the impact of the proposed design choices. Finally, TEncDM outperforms recent diffusion models and some classical autoregressive methods in downstream task experiments.
@@ 10 Limitations
There are three limitations that warrant further investigation. First, the quality of the model can be improved by training diffusion encoder, decoder and denoising model simultaneously. However, we avoid doing so in order to avoid overcomplicating the approach. Second, the samples from the latent space have a high dimensionality that depends on the sequence length, so the training of our method slows down significantly as the length increases. This problem can probably be overcome by training the autoencoder to map a text into a fixed-dimensional latent space, which is a great direction for further research. Third, as different diffusion encoders works better for different tasks, it is necessary to find the best one for each task.
Acknowledgments
This research was supported in part through computational resources of HPC facilities at HSE University.
@@ Appendix B Corruption for decoder training
    Replacing $z_{0}$ with $z_{t}$ by the diffusion forward process, $Cor(z_{0})=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\varepsilon=z_{t}$.
@@ C.2 Wikipedia self-conditioning analysis
Figure 9 depicts the trajectories of prediction magnitudes for varying amounts of generation steps for the Wikipedia dataset. It can be observed that the trajectories exhibit a similar trend to those observed for the ROCStories on Figure 3, although the spread of final prediction magnitudes is larger.
@@ Appendix E Implementation details
Table 6: Hyperparameter values for TEncDM across different datasets.
@@ Appendix G Low MAUVE for RoBERTa encoder
Appendix H Training time comparison
In Table 8 we provide time measurements for one training iteration of the proposed method and its modifications: without diffusion encoder (embedding-based diffusion model) and without self-conditioning. We performed the measurements on a single Nvidia A100 GPU for the batch of 512 texts with a sequence length of 128. We notice that the addition of the diffusion encoder slows down the training by only 15% and the addition of self-conditioning by 20%. Note that the generation speed is the same for models with and without diffusion encoder, as the diffusion encoder is not used during generation.
Model Time for one iteration (s)
TEncDM .$730_{.147}.$
w/o $E_{diff}$$634_{.134}.$
w/o self-conditioning .$605_{.110}.$
Table 8: Training time comparison for TEncDM and its modifications.
Appendix I Scalability
We investigate the scalability of TEncDM. Our model leverages both a text encoder and a diffusion model, offering two avenues for parameter expansion: scaling the diffusion model alone or scaling both the encoder and diffusion model concurrently. In this section, we focus on the first approach, where we exclusively increase the size of the diffusion model. Specifically, we increased the number of transformer heads from 12 to 16, while maintaining the size of each head, and doubled the number of layers from 12 to 24, bringing the total number of parameters to 223M. We call the standard model “base” and the increased model “large”. In Table 9 we show comparison of base and large diffusion models on Wikipedia dataset. The results demonstrate the effectiveness of the scaling strategy. The second approach, involving simultaneous scaling of both the encoder and diffusion model, necessitates significantly greater computational resources and training time. Therefore, we defer the exploration of this approach to future research.
@@ Appendix I Scalability
Appendix J Comparison with GPT-2
@@
Table 11: Comparison with GPT-2 on unconditional generation.
Appendix K Generation examples